# Learning Correlated Astrophysical Foregrounds with Denoising Diffusion Probabilistic Models

### v1

@@ 3.2 Summary Statistics
The first point of comparison is the angular power spectrum, which quantifies scale-dependent correlations in the maps. We compute it using a binned flat-sky power spectrum estimator over a range of $4000300<\ell<$ 4000, in bins of width $\Delta\ell=60$. For the DDPM-generated samples, we initially observe a slight deficit in the power spectrum amplitude, dominant in the Poisson part of the power spectrum. We attribute this to the difficulty of capturing rare, high-intensity pixels—an established limitation of diffusion models [50]. To mitigate this, we apply a post-training rescaling of the pixel intensities. Specifically, we multiply each DDPM sample by the ratio of the standard deviation of all the Agora samples to that of all the generated samples, ensuring the sample variances match. This post-training adjustment proves remarkably effective: it restores agreement with the target power spectrum and also substantially improves the performance of the model across all subsequent summary statistics, without having to retrain the model.
@@ 4.2 Stacked tSZ profiles
Besides the visual inspections mentioned above, we compare the profiles of the tSZ signals from DDPM with Agora. To this end, we use the tSZ simulations and extract $31^{\prime}\times 31^{\prime}$ cutouts centered at the locations of pixels $i$ that fall within a given SNR threshold (A,B) defined as $A\sigma_{T}<|T_{i}-\bar{T}|\leq B\sigma_{T}$, where $\bar{T}$ and $\sigma_{T}$ represent the mean and the standard deviation of the simulated tSZ map. Specifically, we pick regions in the following SNR ranges: (a) $5-10\sigma$, (b) $10-20\sigma$, and (c) $\geq 20\sigma$. Note that the simulated maps are free from instrumental noise, and hence the noise corresponds to the tSZ-confusion noise [46] from the other haloes present in the map. The number of haloes picked from Agora and DDPM simulations agree within 5-10% for the first two cases, but the difference is slightly higher ($\sim 19\%$) for the high SNR case. The larger discrepancy for the high SNR case arises from the difficulty in properly capturing the statistics of the massive clusters with the patch size used for training. We discuss this further in the subsequent sections. In Figure  3, we present the radial profile of the stacked tSZ signals as a function of radius from the center. In the figure, black corresponds to the radial profile from Agora and blue corresponds to DDPM simulations. In the top panels, we also show the difference between Agora and DDPM in yellow, and in the bottom panels, we present the ratio. As evident from the figure, radial profile of the tSZ signals from DDPM agrees with the ones from Agora to within 8%. We also note that the radial profiles from both Agora and DDPM simulations do not reach zero even at $\theta\geq 10^{\prime}$ from the cluster center. This is due to the 2-halo term of the tSZ signal [54] dominating over the 1-halo term at larger distances from the cluster center, and it is being reproduced in DDPM simulations to within 5% when compared to Agora.
@@ 4.3 Power Spectra comparison
Next, we evaluate the auto- and cross- angular power spectra of 200 randomly selected CIB and tSZ patch pairs generated by the DDPM and compare them to patches from the Agora simulation. The CIB exhibits a Poisson-dominated spectrum, where $D_{\ell}$ increases as $\ell^{2}$, consistent with a model with Poissonian shot noise term and a continuous clustering term [55]. In contrast, tSZ shows a peak structure due to the characteristic angular size of galaxy clusters, set by their halo pressure profiles. After applying variance rescaling as described in Section 3.2 to the DDPM outputs, to match the true variance, we find that the generated power spectra match the ground truth within 25% of the sample variance across all multipoles. We find that the CIB power spectrum matches well even without rescaling, while the tSZ spectrum requires rescaling to correct for the underestimation of its extreme-valued pixels. The cross-spectrum between the CIB and tSZ, which probes the spatial correlation between dusty star-forming galaxies and the hot gas traced by tSZ, is also successfully reproduced by the DDPM, well within the expected sample variance.
@@ 4.4 Pixel Intensity Histograms
However, there are discrepancies present in the distribution tails. For the CIB, we observe an underproduction of high-intensity pixels in the DDPM samples, clearly showing a shortfall beyond intensities of $\sim 60\ \mu K$. For the tSZ, the DDPM overestimates the occurrence of strong decrements (large negative values), which correspond to massive clusters. These deviations highlight the challenge in recovering rare, high-amplitude fluctuations that drive the heavy tails of these distributions.
@@ Appendix B Correlations across frequency
To assess the ability of DDPMs to capture the inter-frequency correlations, we trained a model to jointly generate CIB maps at 95, 150, and 857 GHz. The CIB fields at 95 and 150 GHz are expected to be highly correlated, while decorrelation becomes apparent at 857 GHz due to the increased contribution from lower-redshift sources. Figure 8 demonstrates that the DDPM successfully reproduces the correlated structures present in the CIB maps across three frequencies. In addition to visual inspection, we quantitatively assess the model’s ability to reproduce the frequency-dependent correlations using angular power spectra. In Figure 9, we show the relative errors between the DDPM and Agora power spectra for all auto- and cross-power spectra for the three frequencies. The errors are generally within 5%. The right panel shows the cross-correlation coefficients between the 95, 150, and 857 GHz maps. The DDPM captures the near-unity correlation between 95 and 150 GHz and $\sim$ 80% (77%) between 150 (95) GHz and 857 GHz, consistent with the Agora simulations.
@@
Figure 9: Left panel: Relative errors of the DDPM power spectra compared to the Agora test set for all auto- and cross-power spectra of the 95, 150, and 857 GHz CIB maps, averaged over 200 samples. The errors in each spectra is generally within $5\%$. Right Panel: Cross-correlation coefficients between different frequency bands computed from the Agora test maps (solid) and DDPM samples (dashed). The DDPM accurately reproduces the correlations observed in the Agora simulations.

### v3

@@ 3.2 Summary Statistics
The first point of comparison is the angular power spectrum, which quantifies scale-dependent correlations in the maps. We compute it using a binned flat-sky power spectrum estimator over a range of $4000300<\ell<$ 4000, in bins of width $\Delta\ell=60$. For the DDPM-generated samples, we initially observe a slight deficit in the power spectrum amplitude, dominant in the Poisson part of the power spectrum. We attribute this to the difficulty of capturing rare, high-intensity pixels—an established limitation of diffusion models [50]. To mitigate this, we apply a post-training rescaling of the pixel intensities. Specifically, we multiply each DDPM sample by a single global factor: the ratio of the standard deviation of all the Agora samples to that of all the generated samples. For the CIB maps, we find a rescaling factor of $1.0328$, while for the tSZ maps, we obtain $1.1425$. This post-training adjustment proves remarkably effective: it restores agreement with the target power spectrum and also substantially improves the performance of the model across all subsequent summary statistics, without having to retrain the model.
@@ 4.2 Stacked tSZ profiles
Besides the visual inspections mentioned above, we compare the profiles of the tSZ signals from DDPM with Agora. To this end, we use the tSZ simulations and extract $31^{\prime}\times 31^{\prime}$ cutouts centered at the locations of pixels $i$ that fall within a given SNR threshold (A,B) defined as $A\sigma_{T}<|T_{i}-\bar{T}|\leq B\sigma_{T}$, where $\bar{T}$ and $\sigma_{T}$ represent the mean and the standard deviation of the simulated tSZ map. Specifically, we pick regions in the following SNR ranges: (a) $5-10\sigma$, (b) $10-20\sigma$, and (c) $\geq 20\sigma$. Note that the simulated maps are free from instrumental noise, and hence the noise corresponds to the tSZ-confusion noise [46] from the other haloes present in the map. The number of haloes picked from Agora and DDPM simulations agree within 5-10% for the first two cases, but the difference is slightly higher ($\sim 19\%$) for the high SNR case. The larger discrepancy for the high SNR case arises from the difficulty in properly capturing the statistics of the massive clusters with the patch size used for training. We discuss this further in the subsequent sections. Once we collect the cutouts of haloes, we compute the stack in each SNR bin as
$\hat{S}(\theta_{x},\theta_{y})=\frac{\sum_{i}w_{i}S_{i}(\theta_{x},\theta_{y})}{\sum_{i}w_{i}}$ (4.1)
where the index $i$ runs over all the clusters in a given SNR bin defined above and we apply uniform weights $w$ for all the clusters. Here, $\theta_{x}$ and $\theta_{y}$ are Cartesian angular coordinates with the origin at the cluster center. In Figure 3, we present the radial profile of the stacked tSZ signals, $\hat{S}(\theta_{x},\theta_{y})$, as a function of radius $\theta=\sqrt{\theta_{x}^{2}+\theta_{y}^{2}}$ from the cluster center. To compute the radial profile, we average the stacked map $\hat{S}(\theta_{x},\theta_{y})$ over all pixels with coordinates $(\theta_{x},\theta_{y})$ such that $\sqrt{\theta_{x}^{2}+\theta_{y}^{2}}\in(\theta_{b},\theta_{b}+\Delta\theta)$, where $(\theta_{b},\Delta\theta)$ define the edges of the radial bin. We use 10 linearly spaced bins with $\Delta\theta=1^{\prime}$ and $\theta_{\rm max}=10^{\prime}$. In the figure, black corresponds to the radial profile from Agora and blue corresponds to DDPM simulations. In the top panels, we also show the difference between Agora and DDPM in yellow, and in the bottom panels, we present the ratio. As evident from the figure, radial profile of the tSZ signals from DDPM agrees with the ones from Agora to within 8%. We also note that the radial profiles from both Agora and DDPM simulations do not reach zero even at $\theta\geq 10^{\prime}$ from the cluster center. This is due to the 2-halo term of the tSZ signal [54] dominating over the 1-halo term at larger distances from the cluster center, and it is being reproduced in DDPM simulations to within 5% when compared to Agora.
@@ 4.3 Power Spectra comparison
Next, we evaluate the auto- and cross- angular power spectra of 200 randomly selected CIB and tSZ patch pairs generated by the DDPM and compare them to patches from the Agora simulation. The results of this comparison are shown in Figure  4. The CIB exhibits a Poisson-dominated spectrum, where $D_{\ell}$ increases as $\ell^{2}$, consistent with a model with Poissonian shot noise term and a continuous clustering term [55]. In contrast, tSZ shows a peak structure due to the characteristic angular size of galaxy clusters, set by their halo pressure profiles. After applying variance rescaling as described in Section 3.2 to the DDPM outputs, to match the true variance, we find that the generated power spectra match the ground truth within 25% of the sample variance across all multipoles. We find that the CIB power spectrum matches well even without rescaling, while the tSZ spectrum requires rescaling to correct for the underestimation of its extreme-valued pixels. The cross-spectrum between the CIB and tSZ, which probes the spatial correlation between dusty star-forming galaxies and the hot gas traced by tSZ, is also successfully reproduced by the DDPM, well within the expected sample variance.
@@ 4.4 Pixel Intensity Histograms
However, there are discrepancies present in the distribution tails. For the CIB, we observe a slight underproduction of high-intensity pixels, with a mild shortfall beyond intensities of $\sim 60\ \mu K$. For the tSZ, the DDPM underestimates the occurrence of strong decrements (large negative values), which correspond to massive clusters, as seen by the suppressed tails in the unscaled DDPM pixel histogram. These deviations highlight the challenge in recovering rare, high-amplitude fluctuations that drive the heavy tails of these distributions.
@@ 4.6 Bispectrum and Trispectrum
We investigated if the underproduction of extreme-value pixels could be the source of the large-scale discrepancy between Agora and DDPM. To this end, similar to the stacking exercise in § 4.2, we select pixel locations satisfying with ${\rm SNR}\geq 10\sigma$, corresponding to the extreme-value pixels, and use a 2D Gaussian fitting approach to determine the size of the haloes. Next, we mask the locations of these high SNR signals in both Agora and DDPM, and recalculate the bispectrum and trispectrum statistics. While masking these extreme signals reduces the non-Gaussianities suppressing both the bispectrum and the trispectrum, it does not have a noticeable impact on the differences between Agora and DDPM. It is important to note here that we do not retrain the DDPM after masking the extreme-value pixels and leave a more careful investigation of this discrepancy to a future work.
@@ Appendix B Correlations across frequency
To assess the ability of DDPMs to capture the inter-frequency correlations, we trained a model to jointly generate CIB maps at 95, 150, and 857 GHz. The CIB fields at 95 and 150 GHz are expected to be highly correlated, while decorrelation becomes apparent at 857 GHz due to the increased contribution from lower-redshift sources. Figure 8 demonstrates that the DDPM successfully reproduces the correlated structures present in the CIB maps across three frequencies. In addition to visual inspection, we quantitatively assess the model’s ability to reproduce the frequency-dependent correlations using angular power spectra. In Figure 9, we show the standard errors between the DDPM and Agora power spectra for all auto- and cross-power spectra for the three frequencies. The errors are generally within 50% of the standard deviation. The right panel shows the cross-correlation coefficients between the 95, 150, and 857 GHz maps. The DDPM captures the near-unity correlation between 95 and 150 GHz and $\sim$ 80% (77%) between 150 (95) GHz and 857 GHz, consistent with the Agora simulations. We observe that the decorrelation between frequency bands is slightly milder for the DDPM samples at higher multipoles compared to the original Agora simulations. Similar to the investigation done in § 4.6, we mask the extreme-value pixels in the CIB maps and recalculate the correlation coefficients for both Agora and DDPM. The impact of this masking is only marginal and well within the errors shown in Fig. 8. We also defer a careful study of this behavior to a future work.
@@
Figure 9: Left panel: Standard errors of the DDPM power spectra compared to the Agora test set for all auto- and cross-power spectra of the 95, 150, and 857 GHz CIB maps, averaged over 200 samples. The errors in each spectra is generally within $50\%$. Right Panel: Cross-correlation coefficients between different frequency bands computed from the Agora test maps (solid) and DDPM samples (dashed). The DDPM accurately reproduces the correlations observed in the Agora simulations.